OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! Which of the following is equivalent to five and four-ninths divided by seven? Choose one answer. A. seven-ninths B. one and two-sevenths C. five-sevenths D. five and four sixty-thirds

OpenStudy (anonymous):

@happyrosy

OpenStudy (anonymous):

@Jaynator495 @KyanTheDoodle @anonymous97 @eninone

OpenStudy (anonymous):

@hartnn @happyrosy

OpenStudy (teddyiswatshecallsme):

Seven-ninths.

OpenStudy (anonymous):

r u sure?

OpenStudy (teddyiswatshecallsme):

Check it yourself!: http://www.calculatorsoup.com/calculators/math/mixednumbers.php

OpenStudy (anonymous):

ok thx..ur awesome btw

OpenStudy (teddyiswatshecallsme):

You're welcome, and thanks. :D

OpenStudy (anonymous):

np

OpenStudy (teddyiswatshecallsme):

Anymore questions just go ahead and personal message me or tag me. Or both.

OpenStudy (anonymous):

ok i will..thanks

OpenStudy (anonymous):

Jake's band played songs at three different parties. At the first party they played four-fifths as long as they did at the second party, which was one-half as long as the third party. If they played for 6 minutes at the first party, how many minutes did they play at the third party? Choose one answer. A. 2 minutes B. four and four-fifths minutes C. eight and three-fifths minutes D. fifteen minutes

OpenStudy (anonymous):

@Teddyiswatshecallsme i tag you somewhere else and msg

OpenStudy (teddyiswatshecallsme):

Please open this in a new question. :D

OpenStudy (anonymous):

ok

OpenStudy (teddyiswatshecallsme):

Alright. :)

OpenStudy (anonymous):

\[5\frac{ 4 }{ 9 }\div7= \frac{ 9\times5+4 }{ 9}\div7\\ = \frac{ 45+4 }{ 9}\div7\\=\frac{ 49 }{ 9 }\div7\\=\frac{ 49 }{ 9}\times \frac{ 1 }{ 7}\\ =\frac{ 7 }{ 9}\ [seven-ninths]\\ -----LOGICFALL \ Co.-----\]